# Equations of State, Thermodynamics, and Miscibility Curves for Jovian Planet and Giant Exoplanet Evolutionary Models

### v1

@@ 3.3 Schöttler & Redmer (2018)
When the demixing profiles are compared in the abundance of , the abundance for which Morales calculated their miscibility curves, the L0911 and Morales curves are hotter than the SR18 profiles by and K, respectively. Figure 4 shows these profiles (colored lines) compared to adiabats for the SCvH95 and CMS19+HG23 equations of state. If the demixing temperatures of SR18 are used to determine miscibility gaps in gas giant planets, then these will lead to narrower H-He immiscibility regions compared to L0911. The bottom panel of Figure 4 shows decreasing entropy adiabats for both EOSes. Once the adiabat intercepts the miscibility curve, the regions highlighted in red undergo H-He phase separation. For a constant helium fraction of , this occurs at baryon for the CMS19+HG23 EOS and roughly at baryon for the SCvH EOS. We show the L0911 and SR18 miscibility gaps as a function of entropy and helium mass fraction in Figure 5. The gaps grow as entropy and helium mass fraction decrease. As helium mass fraction increases, immiscibility regions open only at lower entropies. Figure 5 illustrates the difference of each miscibility diagram, where L0911 can achieve miscibility gaps at higher entropies than SR18. Since the SR18 curves are significantly cooler, they require lower entropies. Because of this, Mankovich & Fortney (2020) claimed that if helium rain is responsible for the excess heat of Saturn, the demixing temperatures need to be hotter by K. Recently, Brygoo et al. (2021) empirically measured a miscibility curve for , highlighted in red in the top panel of Figure 4. While they found a positive upward shift in the temperature from the SR18 curve, the shift is K, which is far larger than the shift predicted by Mankovich & Fortney (2020).
The SR18 H-He miscibility data have several advantages over the L0911 and Morales data. Its advantage over the L0911 phase diagram is its inclusion of non-ideal entropy of mixing, while both the SR18 and L0911 data have an advantage over the Morales phase diagram in its provision of remixing temperatures for an arbitrary helium mass fraction. Based on this history and due to their inclusion of a dependence on the helium mass fraction, we advocate in the near term the use of the L011 and SR18 curves as the best current options. However, we note that temperature shifts in these curves to alter their demixing temperatures still seem necessary. Such temperature changes echo the suggestions of Nettelmann et al. (2015), Püstow et al. (2016), Mankovich et al. (2016), and Mankovich & Fortney (2020).
@@
We make the L0911 and SR18 immiscibility curves accessible and provide prescriptions to use for giant planet evolutionary and structural calculations. These H-He miscibility prescriptions are allowed to follow the pressure, temperature, and helium abundance of a given planetary interior profile. Therefore, these adapt to the conditions of the interior rather than remaining static; i.e., as the pressure and temperature of the planet interior changes, so does the H-He immiscibility curve. Note that, since detailed evolutionary calculations contain metallicity, these miscibility curves were originally calculated using pure H-He mixtures. Therefore, these curves do not account for alterations of the planetary profiles due to metals. Moreover, the H-He EOSes used to calculate the immiscibility curves by SR18 and L0911 differ from those provided here.
@@ 4 Thermodynamic Consistency
In Figure 6, we show regions of thermodynamic space relevant to planetary evolution with shaded regions where each EOS deviates from consistency. The SCvH95 EOS (top row) demonstrates more thermodynamic consistency in important regions for gas giant interiors ( g cm), showing consistency in these regions. The CMS19+HG23 EOS (bottom row) shows thermodynamic consistency in these regions, in approximate agreement with Becker et al. (2014) in their thermodynamic consistency tests. Thermodynamic inconsistencies imply that energy conservation of planetary evolution models can only be as good as their EOSes allow.
5 Generalized Convection Criteria and The Energy Equation
In the standard mixing-length convection theory (Weiss et al., 2004; Kippenhahn et al., 2012; Joyce & Tayar, 2023), the associated gedanken thought experiment, a fluid element is displaced and kept isolated at constant pressure with its new surroundings. Isolation in this context means the fluid element maintains a constant entropy and composition. Various structural and thermodynamic derivatives emerge to be important in the context of such theories of convection and convective transport. For simplicity, we evaluate these criteria assuming that the only compositional variation is in the helium mass fraction, , but these expressions can easily be generalized to arbitrarily complex compositional variation. The derivatives in question include ( is the pressure scale height):
@@ 5 Generalized Convection Criteria and The Energy Equation
An exhaustive table of the available values and derivatives for each H-He EOSes is shown in Table 1. Units used in our code are listed in Table 2.
Note. — These quantities and derivatives are also available as a function of . The logarithmic quantites are log base 10, and the units are tabulated in Table 2. The original CMS19 tables were provided by their authors in the bases and . To ensure consistency, we used the original pressure-temperature combined with the HG23 components to obtain new tables. All quantities and derivatives of the CMS19 and MLS22 tables contain these non-ideal corrections. The thermodynamic quantities and derivatives are also available for pure water (AQUA), post-perovskite, serpentine, and iron, along with the provided “metal” mixture tables.
@@ 5.1 Convection
The relative density of this isolated fluid compared to that of its surroundings is the criterion for convective stability. In the absence of a composition gradient, the Schwarzchild criterion for convective stability is
@@
such that . The split into and is particularly useful. To whit, the superadiabaticity, , can be shown to be
so the Schwarzchild criterion for convective stability can then be expressed as
It is important to note that, in the presence of a composition gradient, marginally Schwarzchild stable regions are not necessarily isentropic. An isentropic structure must meet the criterion
@@
In cases where a fluid is stable against convection according to the Ledoux criterion, but Schwarzchild unstable, semiconvection can arise in the form of double diffusive convection (see Walin, 1964; Kato, 1966; Stevenson, 1985; Rosenblum et al., 2011; Leconte & Chabrier, 2012; Mirouh et al., 2012; Wood et al., 2013; Leconte & Chabrier, 2013; Nettelmann et al., 2015; Moll et al., 2016; Garaud, 2018, and references therein). In lieu of any detailed understanding of energy transport in semiconvective regions (Debras et al., 2021b; Mankovich & Fortney, 2020; Mankovich et al., 2016; Nettelmann et al., 2015; Leconte & Chabrier, 2012), planetary models have included a density ratio , which describes the degree to which temperature and composition stratifications contribute to the density stratification to trigger convection. This parameter is defined as (see Rosenblum et al., 2011; Mirouh et al., 2012, for discussions)
where represents the overall composition gradient. Eq. 23 can be written with as
@@
The parameter can be taken as a free parameter that stretches from 0 (Schwarzchild) to 1 (Ledoux). In terms of entropy, the criterion becomes
If , then Eq. 35 reduces to the Schwarzchild criterion, Eq. 29. If , then Eq. 35 reduces to the Ledoux criterion, Eq. 32, modulo the factor.
@@
In the repository, we provide these derivatives in functional and tabulated form, with the choice of parameter . Figure 7 shows the sign changes of the relevant derivatives for the convection criteria at typical values in the interior of gas giants. A negative ensures that when the Schwarzchild convection stability condition is violated (i.e., convection occurs), the resulting entropy gradient is positive, assuming that the composition gradient () is negative. This would also hold for the Ledoux condition against convection but for . The sign change of this derivative at interior conditions indicates that the reverse occurs; namely, the entropy profile gradient becomes negative when the Ledoux convection stability condition is violated in these regions.
@@
We also provide the expressions to compute the Brunt-Väisälä frequency (), which is the bouyant oscillation frequency in a stable stratified fluid. This frequency is defined as (Lattimer & Mazurek, 1981; Unno et al., 1989; Saio, 1993; Passamonti et al., 2009; Fuller et al., 2016; Mankovich & Fuller, 2021, also shown in Appendix B.):
which can be shown444in Appendix B to be equal to
@@ 6 Conclusion
With this paper, and its associated website (https://github.com/Rob685/hhe_eos_misc), we provide a more centralized and organized way to access the crucial hydrogen-helium mixture equations of state, accompanied by the various derivatives necessary to calculate heat transport and different convection criteria. The EOSes and derivatives are recalculated in the range Mbar and Kelvin, a range that is comparable to each of the EOS tables that we have adapted. This work allows for the prompt calculation of the derivatives required in the convection criteria for stability, namely, the partial derivatives in Eqs. 35 and 37. Importantly, we calculate the pseudo-chemical potential (Eq. 40) so that it is readily available for energy calculations for any helium mixture.
Furthermore, we make the miscibility curves from L0911 and SR18 accessible to the community of planetary modelers. As discussed in Section 3.3, we provide the ability to calculate the miscibility curves for general planetary profiles without restricting to a single value of . The present work marks a step toward addressing the many inconsistencies between EOSes, miscibility curves, and other physics relevant to planetary evolution. Important future work by the community includes the calculation of self-consistent H-He immiscibility and EOS curves that incorporate all nonideal interactions between H-He and metals in a self-consistent fashion. Such a future project will remove many of the remaining ambiguities in the thermodynamics of the giant planets of our solar system and beyond.

### v2

@@ 3.3 Schöttler & Redmer (2018)
When the demixing profiles are compared in the abundance of 444This is the primordial solar helium abundance value obtained from from Bahcall et al. (2006), where , ., the abundance for which Morales calculated their miscibility curves, the L0911 and Morales curves are hotter than the SR18 profiles by K and K, respectively. The top panel of Figure 4 shows these demixing temperature profiles (colored lines) compared to adiabats for the SCvH95 and CMS19+HG23 equations of state. If the demixing temperatures of SR18 are used to determine miscibility gaps in gas giant planets, then these will lead to narrower H-He immiscibility regions compared to L0911. The bottom panel of Figure 4 shows decreasing entropy adiabats for both EOSes. Once the adiabat intercepts the miscibility curve, the regions highlighted in red undergo H-He phase separation. For a constant helium fraction of , this occurs at baryon for the CMS19+HG23 EOS and roughly at baryon for the SCvH EOS. We show the L0911 and SR18 miscibility gaps as a function of entropy and helium mass fraction in Figure 5. The helium rain regions grow as entropy and helium mass fraction decrease. As helium mass fraction increases, immiscibility regions open only at lower entropies. Figure 5 illustrates the difference of each miscibility diagram, where L0911 can achieve miscibility gaps at higher entropies than SR18. Since the SR18 curves are significantly cooler, they require lower entropies. Because of this, Mankovich & Fortney (2020) found that if helium rain is responsible for the excess heat of Saturn, the H-He demixing temperatures of the SR18 miscibility curves need to be shifted by K. Recently, Brygoo et al. (2021) empirically measured a miscibility curve for , highlighted in red in the top panel of Figure 4. While they found a positive upward shift in the temperature from the SR18 curve, the shift is K, which is far larger than the shift predicted by Mankovich & Fortney (2020).
@@
The SR18 H-He miscibility data have several advantages over the L0911 and Morales data. Its advantage over the L0911 phase diagram is its inclusion of non-ideal entropy of mixing, while both the SR18 and L0911 data have an advantage over the Morales phase diagram in its provision of demixing temperatures for an arbitrary helium mass fraction. We make the L0911 and SR18, and the Brygoo et al. (2021) experimental immiscibility curves conveniently accessible to use for giant planet evolutionary and structural calculations555All smooth miscibility curves are provided here: https://github.com/Rob685/hhe_eos_misc/tree/main/misc. These H-He miscibility prescriptions are allowed to follow the pressure, temperature, and helium abundance of a given planetary interior profile. These adapt to the conditions of the interior rather than remaining static; i.e., as the pressure and temperature of the planet interior changes, so does the H-He immiscibility curve. Note that these miscibility curves were originally calculated using pure H-He mixtures. Therefore, these curves do not account for alterations of the planetary profiles due to metals. Moreover, the H-He EOSes used to calculate the immiscibility curves by SR18 and L0911 differ from those provided here.
4 EOS & Miscibility tables and methods
This work provides three additional inverted tables in addition to the original tables published by SCvH95, MH13, CMS19, CD21, and MLS22. As shown in Table 1, the tables computed here have the following independent thermodynamic variable pairs: ; ; and . These tables have been computed across values of and . All table ranges are described below. While SCvH95 and CMS19 have published tables in both pressure-temperature and density-temperature, EOS mixtures of and using the AVL (Eqs. 5 and 6) can be performed only when combining entropy, density, and internal energy (e.g., Saumon et al., 1995). Tables with as independent variables require this density be the density of the mixture and the temperature of the mixture. Likewise, tables with independent variables require that the particular independent entropy and density be the mixture quantities, and not those of the original H and He tables.
To meet this requirement, numerical inversions are performed to obtain the appropriate independent thermodynamic variables; i.e., if an EOS call requires the temperature as a function of entropy, pressure, Y, and Z, then the entropy and pressure should be that of the H-He-metal mixture, and the appropriate temperature should be returned. We use a modified Powell method from MINPACK (More et al., 1980) as applied in SciPy (Virtanen et al., 2020)666The multi-variable inversions were performed using Scipy’s root application of MINPACK’s hybr FORTRAN routines. using an ideal H-He EOS as an initial guess. Such root finding methods allows us to recover original table values, and values between the original table points, to fractional errors between to . All original and subsequent inverted tables are interpolated using piecewise multi-linear methods (Weiser & Zarantonello, 1988; Virtanen et al., 2020).777The 4-dimensional linear interpolations are performed via Scipy.
4.1 EOS table ranges
We list all the thermodynamic variables for each table in Table 1 and the units are specified in Table 2. Each H-He EOS is combined with the pure water AQUA table (Haldemann et al., 2020). For brevity, each H-He-Z mixture table name is abbreviated with +AQUA to indicate a H-He-Z mixture. The () CMS19+HG23+AQUA and CD21+AQUA tables range from kbaryon to kbaryon, and the SCvH95+AQUA and MLS22+HG23+AQUA () tables range from kbaryon to kbaryon. All tables range from to and from to . The () tables cover the original temperature ranges, except for temperatures above K. 888All the EOS tables can be extrapolated linearly from the bounds indicated, but extrapolation is not advised. The H-He EOS tables are square in shape and are designed to cover the thermodynamic space of gas-giant planet evolution, but are not yet designed to cover temperatures above Kelvin and densities above 100 g cm-3.
4.2 H-He and Miscibility Package description
Each miscibility curve has its own module:
L0911: misc_lor.py
SR18: misc_sch.py
Brygoo et al. (2021): misc_brygoo.py .
Each of these modules contains T in units of K and P in Mbar, where Tm is the demixing temperature of each curve as a function of pressure and helium mass fraction and Pm is its inversion. Each of these functions has an option to offset the demixing temperatures by an amount in Kelvin. These functions were used to compute and show the miscibility curves in Figures 3 and 4.
A wrapper module called misc.py calls each of these files and offers users a method to calculate intercepting points between a choice of a miscibility curve and any temperature-pressure profile to generate a miscibility region. This function is used in APPLE’s miscibility scheme 1 (Sur et al., 2024). The misc.py module also contains a way to calculate Y, or the minimum helium abundance required for H-He immiscibility for a given pressure and temperature, as well as its derivative with respect to temperature. The Ym function is used in APPLE’s miscibility scheme 2.
As with the miscibility curves, and as described in the EOS repository tutorials999https://github.com/Rob685/hhe_eos_misc/blob/main/eos_tutorial.ipynb, each H-He EOS shown in Figure 1 has its own EOS module:
SCvH95: scvh_eos.py
MH13: mh13_eos.py
CMS19 (with or without HG23): cms_eos.py
CD21: cd_eos.py
MLS22: mls_eos.py
as well as the Z tables:
Water (AQUA): aqua_eos.py
Post-perovskite: ppv_eos.py
Iron: fe_eos.py
Serpentine (ANEOS): serpentine_eos.py .
For H-He-Z mixtures, we include a module called mixtures_eos.py that combines the user’s choice of H-He EOS with the user’s choice of Z EOS. The post-perovskite and iron EOS tables are provided by Zhang & Rogers (2022) via private communication, and the serpentine EOS from the analytical equation of state (ANEOS Thompson, 1990). All of these files contain the pre-calculated inverted tables described in this section and in Table 1, the inversion functions used for each table, and the derivatives.
4.3 Derivatives
All EOS modules described and listed above offer their own derivatives, but, as described in our repository derivatives tutorial101010https://github.com/Rob685/hhe_eos_misc/blob/main/eos_derivatives_tutorial.ipynb, mixtures_eos.py houses all derivatives in Table 1 for each H-He-Z mixture. For each derivative, the respective tables shown in the column headers of Table 1 are called to perform either a first order or second order finite difference. The user reserves control over the finite-difference deltas. Examples of a few derivatives used in energy transport are shown in Figure 7 and Figure 8 using fractional differences of along constant entropies and pressures for the purpose of differentiation. The derivatives shown in Figure 8 are used in APPLE to compute the Schwarzchild criteria () and the Ledoux criteria ). We discuss our convection convention in Section 6. Each of the files listed in this section can be accessed by following the installation instructions in the repository page.
@@ 5 Thermodynamic Consistency
In Figure 6, we show regions of thermodynamic space relevant to planetary evolution with shaded regions where each EOS deviates from consistency. The top row shows the thermodynamic consistency of the CMS19+HG23 EOS, the middle row that of the CD21 EOS, while the bottom row shows the same for the SCvH95 EOS. Thermodynamic consistency is defined here as the difference between Equations 9 and 10 and the EOS table values over a 2-D grid of entropy and density values. In the CMS19+HG23 EOS, the metallic hydrogen phase transition is less thermodynamically consistent (denoted by darker colors), with overall errors of in the temperature. In the SCvH95 EOS, less thermodynamically consistent regions are found at higher entropies, due to interpolation techniques used by SCvH95 across the molecular/metallic transition and the plasma phase transition, with overall errors of in the temperature. These inconsistencies are in approximate agreement with the thermodynamic (in)consistency seen by Becker et al. (2014), and they imply that energy conservation of planetary evolution models can only be as good as their EOSes allow. Crucially, the internal energy of the CMS19+HG23 EOS does not yet account for non-ideal interactions. This inconsistency should be partially resolved once these internal energy terms become available. On the other hand, the internal energy values of the CD21 EOS already account for these non-ideal interactions, albeit at a constant helium fraction. Unlike the SCvH95 EOS, the CD21 does not interpolate across the plasma phase transition region. This accounts for the thermodynamic consistency observed in the CD21 EOS of under . Thermodynamic consistency must be considered when accounting for energy transport and overall energy conservation in evolutionary calculations.
6 Convection Criteria and The Energy Equation
In the standard mixing-length convection theory (Weiss et al., 2004; Kippenhahn & Weigert, 1990; Kippenhahn et al., 2012; Joyce & Tayar, 2023), the associated gedanken thought experiment, a fluid element is displaced and kept isolated at constant pressure with its new surroundings. Isolation in this context means the fluid element maintains a constant entropy and composition. Various structural and thermodynamic derivatives emerge to be important in the context of such theories of convection and convective transport. For simplicity, we evaluate these criteria assuming that the only compositional variation is in the helium mass fraction, , but these expressions can easily be generalized to arbitrarily complex compositional variation. The derivatives in question include:
@@ 6 Convection Criteria and The Energy Equation
An exhaustive table of the available values and derivatives for each H-He EOSes is shown in Table 1. Units used in our code are listed in Table 2. In Figure 7, we present , , , and . The entropy values selected for and encompass a range relevant to gas giant planet evolution, and the pressures selected for , and encompass the interior regions relevant for H-He phase separation.
Note. — The thermodynamic quantities and derivatives are also available for pure water (AQUA), post-perovskite, serpentine, and iron, along with the provided “metal” mixture tables111111All tables are accessible at our EOS repository: https://github.com/Rob685/hhe_eos_misc. The logarithmic quantites are log base 10, and the units are tabulated in Table 2. The original CMS19 tables were provided by their authors in the bases and . To ensure consistency, we used the original pressure-temperature combined with the HG23 components to obtain new tables. All quantities and derivatives of the CMS19 and MLS22 tables contain these non-ideal corrections.
@@ 6.1 Convection
The relative density of this isolated fluid compared to that of its surroundings is the criterion for convective stability. In the absence of a composition gradient, the Schwarzschild criterion for convective stability is
@@
such that and is the pressure scale height, . The split into and is particularly useful. To whit, the superadiabaticity, , can be shown to be
so the Schwarzschild criterion for convective stability can then be expressed as
Recall that in the absence of a composition gradient, marginal Schwarzschild stability (i.e., the transition from stability to instability) corresponds to isentropy (). However, Eq. 28 shows that in the presence of a composition gradient, the transition from stability to instability occurs for a non-zero entropy gradient (even if ).
@@
In cases where a fluid is stable against convection according to the Ledoux criterion, but Schwarzschild unstable, semiconvection can arise in the form of double diffusive convection (see Walin, 1964; Kato, 1966; Stevenson, 1985; Rosenblum et al., 2011; Leconte & Chabrier, 2012; Mirouh et al., 2012; Wood et al., 2013; Leconte & Chabrier, 2013; Nettelmann et al., 2015; Moll et al., 2016; Garaud, 2018, and references therein). In lieu of any detailed understanding of energy transport in semiconvective regions (Debras et al., 2021b; Mankovich & Fortney, 2020; Mankovich et al., 2016; Nettelmann et al., 2015; Leconte & Chabrier, 2012), recent planetary models have included a density ratio , which describes the degree to which temperature and composition stratifications contribute to the density stratification to trigger convection. This parameter is defined as (see Rosenblum et al., 2011; Mirouh et al., 2012, for discussions)
Eq. 22 can be written with as
@@
The parameter can be taken as a free parameter that stretches from 0 (Schwarzschild) to 1 (Ledoux). In terms of entropy, the criterion becomes
If , then Eq. 33 reduces to the Schwarzschild criterion, Eq. 28. If , then Eq. 33 reduces to the Ledoux criterion, Eq. 30, modulo the factor.
@@
In the repository, we provide these derivatives in functional form. Figure 8 shows the sign changes of the relevant derivatives for the convection criteria at typical values in the interior of gas giants. Note that when the composition gradient () is negative, a negative ensures that when the Schwarzschild convection stability condition is violated, the resulting entropy gradient is positive. This also holds for the Ledoux condition against convection, but for instead of . A sign change (see Figure 8) of at interior conditions indicates that the reverse occurs; namely, the entropy profile gradient becomes negative when the Ledoux convection stability condition is violated in these regions. The derivative at constant () increases with decreasing entropy for densities g cm-3. Note also that temperatures here and for realistic gas giants are significantly higher than the hydrogen solidus line found in the CMS19 EOS. The SCvH95 EOS (dashed lines) also exhibits this behavior. This is in contrast with the Schwarzchild derivative, , which has no sign change.
@@
We also provide the expressions to compute the Brunt-Väisälä frequency (), which is the bouyant oscillation frequency in a stable stratified fluid. This frequency is defined as (Lattimer & Mazurek, 1981; Unno et al., 1989; Saio, 1993; Passamonti et al., 2009; Fuller et al., 2016; Mankovich & Fuller, 2021, also shown in Appendix A.):
which can be shown121212in Appendix A to be equal to
6.2 Latent Heat Correction Due to Helium Phase Separation
We note that the energy source that is the gravitational settling of helium should exceed by a large margin the energy source that is the latent heat of helium phase separation and that all current evolutionary codes properly handle the former. Though we emphasize that the Ledoux condition (Eqs 22 and 30) assumes that the fluid mixtures are fully miscible, something not expected in a multi-phase context, one can approximately generalize the Ledoux condition to account for latent heat evolution (though in a miscible context) by employing what is done in the meteorological literature (Stevens, 2005) vis à vis wet versus dry adiabats and by employing the generalization found in Tremblin et al. (2019). The result is a multiplication factor on (Eqs. 21 and 22) of the form:
where is the latent heat per mass and Y is the helium abundance at which immiscibility ensues at a given and . Assuming the latent heat estimate of Stevenson & Salpeter (1977b) of order to per helium nucleus, the magnitude of this correction at the immiscibility curve is generally close to one, ranging from 0.93 to 0.99 for a broad range of temperature/pressure points and miscibility models. For the helium rain regions shown in Figure 9, it is indeed close to one. Above the immiscibility curve this correction is one and far below it one must have a better sense of the variation in the latent heat as the helium abundance associated with the phase partitioning varies. However, until complete numerical data on the latent heats under all thermodynamic and compositional circumstances are available, the helium phase-separation latent heat effect on a generalized Ledoux criterion must, for all intents and purposes, be considered unknown.
@@ 6.3 The energy equation
7 One Jupiter Mass model examples
As an example of the use of these EOS tables, we employ the evolutionary code APPLE (Sur et al., 2024) to compute a pair of one MJ evolutionary models, with and without helium rain. We assume here the CD21 equation of state and the Lorenzen et al. (2011) miscibility curves, shifted by K131313Although this model is not designed to fit Jupiter in detail, such a shift is required to maintain a flat 3 solar metallicity envelope to closely match the observed abundance (Von Zahn et al., 1998).. These models do not purport to be fits to Jupiter, as a full and detailed study with APPLE is the subject of upcoming work (Sur et al. 2024b, in prep). The APPLE code includes, as best we currently can, helium rain. In it, we employ a scale parameter (, here set equal to cm) that governs the magnitude of an settling/advective term, while conserving mass and properly accounting for the gravitational energy release. The parameter results in an equilibrium helium “rain cloud” that is quasi-exponentially distributed in space (similar to models of condensation clouds, but consistent with the miscibility curves).
Figure 9 shows the thermal profile evolution as a function of pressure and mass on the left column, and the evolution of the entropy profiles and helium profiles on the right column. These models were computed using a flat water abundance profile at 3 times solar throughout the envelope, with a total metallicity mass of M⊕ for both models, assuming M⊕ of ice/rock/iron mixtures in a compact core. We used atmospheric boundary conditions that included solar irradiation and ammonia clouds as computed by Chen et al. (2023). While these models are merely test examples of the use of our tables in an evolutionary code, they do still incorporate much of the basic physics necessary to address gas giant evolution. More sophisticated models using APPLE will be forthcoming.
@@ 8 Conclusion
With this paper, and its associated repository, we provide a more centralized and organized way to access the crucial hydrogen-helium-metallicity mixture equations of state, accompanied by the various derivatives necessary to calculate heat transport and different convection criteria. The EOSes and derivatives are recalculated in the range and , a range that is comparable to each of the EOS tables that we have adapted. Importantly, we calculate the pseudo-chemical potential (Eq. 39) so that it is readily available for energy calculations for any helium mixture.
Our primary purpose is to provide curated access to equation of state and miscibility data and thermodynamic derivatives presently scattered in the literature. In addition, though the standard Schwarzchild and Ledoux conditions discussed in the stellar and planetary evolution literature (Kippenhahn & Weigert, 1990; Paxton et al., 2011; Gabriel et al., 2014; Mihalas, 1978) use temperature and composition as primary variables and refer centrally to and , the alternate set of variables (entropy and composition) have virtues (see Sur et al., 2024) and in this paper have provided the corresponding formulae.
The reader should note that the effects of radiative/conductive and composition diffusion on the Ledoux condition and the various approaches to doubly-diffusive instabilities (“semi-convection”) are discussed in many works (see, e.g., Kippenhahn & Weigert (1990), Paxton et al. (2011), Langer et al. (1983), and Tremblin et al. (2019)) and will not be repeated here. On doubly-diffusive instabilities, the group of Pascale Garaud has generated a series of important papers (Rosenblum et al., 2011; Traxler et al., 2011; Mirouh et al., 2012; Wood et al., 2013; Brown et al., 2013; Moll et al., 2016; Tulekeyev et al., 2024) in the giant planet context. While a full discussion of these additional issues is beyond the scope of this paper, mention should be made of the complicating effects of helium rain. The standard Ledoux condition doesn’t explicitly incorporate the latent heat of miscibility, nor the issue of the growth and advective settling of helium droplets in a gravitational field. Importantly, the latent heat of hydrogen/helium immiscibility for the various miscibility models is not provided in the papers that perform detailed miscibility calculations. As a result, it is premature to delve into this topic in detail until such data from modern calculations are available. We note that the latent heat effect is likely much smaller than the heating per helium nucleus provided by the settling rain itself in a gravitational field, and the latter effect is properly handled in extant giant planet evolution codes (e.g., Sur et al., 2024).
In addition, a study of the instantaneous effect of helium droplet settling on a generalized Ledoux condition must acknowledge that the regions of the giant planet where miscibility emerges are likely already convective and that they cool into the miscibility curve. This implies that there is already upwelling and mixing in the region for which there is incipient phase separation. This hydrodynamics is not well addressed by the standard linear stability gedanken experiment that leads to the Ledoux condition, or its standard generalizations. Moreover, mixing-length theory, the standard context of spherical evolutionary codes, is not well suited to this multi-phase, multi-dimensional physics. Clearly, this highlights the need for a more sophisticated approach going forward to the interesting physics of helium rain in planet evolutionary models.
In summary, this work is designed to facilitate the prompt calculation of the thermodynamic quantities and their derivatives required for updated giant planet and exoplanet modeling in the Juno and Cassini eras and to help new groups of modelers join this exciting field.
Data Availability
All data are available on Zenodo under an open-source Creative Commons Attribution license: https://doi.org/10.5281/zenodo.10659249 (catalog doi:10.5281/zenodo.10659249).